# On the constrained feedback linearization control based on the MILP representation of a ReLU-ANN

### v2

@@ IV-B Horizontal 1D quadrotor (Quad-1D)
$\Phi(\boldsymbol{z},v)=\tau(v+\gamma z_{3})/\left(\Gamma\sqrt{1-\zeta^{2}(\boldsymbol{z})}\right)+\sin^{-1}\zeta(\boldsymbol{z}),$ (24)
@@
Figure 5: The input $u=\Phi(\cdot)$, the approximation $\Phi_{NN}(\cdot)$ and its predicted trajectories.$|u|=\pm{u}_{max}$ (blue) and $u=\pm({u}_{max}-\epsilon)$ (green).
To construct the MPC (21), the linearized system is rewritten in the form of (3) and discretized with the forth-order Runge-Kutta method, sampled at $t_{s}=100$ms. The prediction horizon is $N_{p}=10$ steps, the time-varying reference $\boldsymbol{z}_{ref}$ is depicted in gray on Fig. 6 and the cost function $\ell(\cdot)$ as in (21a) is chosen as: $\ell(\boldsymbol{z},v)=\|\boldsymbol{z}-\boldsymbol{z}_{ref}\|_{\boldsymbol{Q}}^{2}+\|v\|_{\boldsymbol{R}}^{2},$ with $\boldsymbol{Q}=\begin{bmatrix}10&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$, $\boldsymbol{R}=0.0175$, $\boldsymbol{z}(0)=\begin{bmatrix}-0.5&-0.15&-0.2\end{bmatrix}^{\top}$. The system’s trajectory is given in Fig. 5 and 6 while the computational effort in both examples is reported in Table II with the min, max and mean value of the calculation time spent in the control loops.
Figure 6: Trajectory tracking with the MPC (21) for (23).
TABLE II: Computation time of the simulations (ms).
@@ IV-C Discussion
For system (22), from Table II, it is noteworthy that both Lcost and Qcost controllers can be implemented with relatively low computation time (from 20-40ms). The only shortcoming is the discontinuity of the Lcost controller, which is caused by the non-smooth characteristics of the constraints (15b)–(15d) in combination with the linear cost. Yet, the problem is not present for the Qcost controller, and interestingly, the computation cost is even lower as compared to that of Lcost. Stable tracking can also be found in the example of the 1D quadrotor (23), where a time-varying trajectory can be tracked with the advantage of anticipation from MPC (see Fig. 6). Given the error bound as in (8), the input constraint is guaranteed by imposing $|\Phi_{NN}(\cdot)|\leq{u}_{max}-\epsilon$. It is crucial to emphasize that in our approach, neither the predicted constraint nor the model underwent approximation through state feedback, as observed in [13, 11]. Hence, the constraints were rigorously enforced over the prediction horizon for the nominal dynamics. This advancement, from a theoretical standpoint, paves the way for constrained FL-based control design with rigorously guaranteed stability. One potential resolution involves adopting the standard MPC axioms [26] incorporating a stabilizing selection of parameters $N_{p},\mathcal{X}_{z}$ and $\ell(\cdot)$ in (21). This reiterates the necessity of delving deeper into understanding the proposed MI linear constraint (9) on a case-by-case basis.
@@
Finally, although satisfactory results and general development were found, the real-time implementation with MPC is evidently still hindered by the solving of the MIQP (21) (see Fig. 7), especially when the number of binary variables grows with the number of neurons and the prediction horizon222 Each time (21b) is imposed, $\sum_{k=1}^{K-1}n_{k}$ binary variables are used. Hence, over $N_{p}$ steps, the number of binary variables is $N_{p}\textstyle\sum_{k=1}^{K-1}n_{k}$. $\left(N_{p}\textstyle\sum_{k=1}^{K-1}n_{k}\right)$. This implementational shortcoming becomes even more evident when multi-input systems are considered, since the linearizing mapping will certainly require a larger network. For the method to be real-time capable, future work concerns the simplification of the constraints and an optimal choice of the ReLU-ANN parameters.
V Conclusion and outlook
@@
This work addressed the feedback linearization control problem with the corresponding convoluted constraints propagated through the linearizing coordinate change. The constraint characterization includes approximating the constraints by a rectified linear unit artificial neural network and reformulating it into the equivalent mixed-integer linear constraints. The applicability and validity of the proposed method were shown via simulation tests with the control Lyapunov function-control barrier function, and model predictive control strategies. Future theoretical directions focus on the investigation of the new constraints and their adaptation within the standard control design for linear systems, showcasing the advantage with comparative study. Meanwhile, practical extension will target optimization in the design parameters and the method’s applications to multi-input systems.

### v3

@@ 4.2 Horizontal 1D quadrotor (Quad-1D)
$\Phi(\boldsymbol{z},v)=\tau(v+\gamma z_{3})/(\Gamma\sqrt{1-\zeta^{2}(\boldsymbol{z})})+\sin^{-1}\zeta(\boldsymbol{z}),$ (24)
@@
Figure 6: The input $u=\Phi(\cdot)$, the approximation $\Phi_{NN}(\cdot)$ and its predicted trajectories.$|u|=\pm{u}_{max}$ (blue) and $u=\pm{u}_{max}^{\epsilon}$ (green).
To construct the MPC (21), the linearized system is rewritten in the form of (3) and discretized with the forth-order Runge-Kutta method, sampled at $t_{s}=100$ms. The prediction horizon is $N_{p}=10$ steps, the time-varying reference $\boldsymbol{z}_{ref}$ is depicted in gray on Fig. 5 and the cost function $\ell(\cdot)$ as in (21a) is chosen as: $\ell(\boldsymbol{z},v)=\|\boldsymbol{z}-\boldsymbol{z}_{ref}\|_{\boldsymbol{Q}}^{2}+\|v\|_{\boldsymbol{R}}^{2},$ with $\boldsymbol{Q}=\begin{bmatrix}10&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$, $\boldsymbol{R}=0.0175$, $\boldsymbol{z}(0)=\begin{bmatrix}-0.5&-0.15&-0.2\end{bmatrix}^{\top}$. The system’s trajectory is given in Fig. 6 and 5, while the online computation time in both examples is reported in Table 2 with its min, max and mean value.
Table 2: Computation time of the simulations (ms).
@@ 4.3 Discussion
For system (22), from Table 2, it is noteworthy that both Lcost and Qcost controllers can be implemented with relatively low computation time (from 20-40ms). The only shortcoming is the discontinuity of the Lcost controller, which is caused by the non-smooth characteristics of the constraints (15b)–(15d) in combination with the linear cost. Yet, the problem is not present for the Qcost controller, and interestingly, the computation cost is even lower as compared to that of Lcost. Stable tracking can also be found in the example of the 1D quadrotor (23), where a time-varying trajectory can be tracked with the advantage of anticipation from MPC (see Fig. 5). Given the error bound as in (8), the input constraint is guaranteed by imposing $|\Phi_{NN}(\cdot)|\leq{u}_{max}^{\epsilon}$. It is crucial to emphasize that in our approach, neither the predicted constraint nor the model underwent approximation through state feedback, as observed in [13, 11]. Hence, the constraints were rigorously enforced over the prediction horizon for the nominal dynamics. This advancement, from a theoretical standpoint, paves the way for constrained FL-based control design with rigorously guaranteed stability. One potential resolution involves adopting the standard MPC axioms [26], incorporating a stabilizing selection of parameters $N_{p},\mathcal{X}_{z}$ and $\ell(\cdot)$ in (21). This reiterates the necessity of delving deeper into understanding the proposed MI linear constraint (9) on a case-by-case basis.
@@
Finally, although satisfactory results and general development were found, the real-time implementation with MPC is evidently still hindered by the solving of the MIQP (21) (see Fig. 7), especially when the number of binary variables grows with the number of neurons and the prediction horizon222 Each time (21b) is imposed, $\sum_{k=1}^{K-1}n_{k}$ binary variables are used. Hence, over $N_{p}$ steps, the number of binary variables is $N_{p}\textstyle\sum_{k=1}^{K-1}n_{k}$. $\left(N_{p}\textstyle\sum_{k=1}^{K-1}n_{k}\right)$. This implementational shortcoming becomes even more evident when multi-input systems are considered, since the linearizing mapping will certainly require a larger network. For the method to be real-time capable, future work concerns the simplification of the constraints and an optimal choice of the network parameters.
5 Conclusion and outlook
@@
This work addressed the feedback linearization control problem with the convoluted constraints caused by the linearizing coordinate change. The constraint characterization includes approximating the mapping by a rectified linear unit artificial neural network and reformulating it into the equivalent mixed-integer linear constraints. The applicability and validity of the proposed method were shown via simulation tests with the control Lyapunov function-control barrier function, and model predictive control strategies. Future theoretical directions focus on the investigation of the new constraints and their adaptation within the standard linear control design, showcasing the advantage with comparative study. Meanwhile, practical extensions will target optimization in the design parameters and the method’s applications to multi-input systems.